OpenStudy (anonymous):

The net price of a refrigerator is $1500. If the refrigerator has a trade discount of 20%, find the list price of the refrigerator.

OpenStudy (tennistar):

Lrf - .2Lrf = Net price Lrf (1-.2) = 1500 Lrf = 1500 / 0.8 list price was 1875.

OpenStudy (anonymous):

actual price=$1500 discount=(20/100)*1500 so the final price =(actual price)-(discount)

OpenStudy (anonymous):

great thanks, that was a question on my quiz and I couldnt figure out how they got the answer.